Find limit of
f(x,y)=(x^3y^3)/(x^3+y^3) at (x,y)=(0,0)
1
Expert's answer
2021-08-22T17:48:10-0400

"lim_{(x,y)\to (0,0)} \frac{x^3y^3}{x^3+y^3}\\
\text{put}y=mx\\
lim_{x\to 0} \frac{x^3}{1+m^3}\\
=0\\
\text{Thus, limit exists and is equal to 0.}"